# Dual-Stream Diffusion for World-Model Augmented Vision-Language-Action Model

### v1

@@ 4.3 Vision-action joint sampling and inference-time scaling
During inference, we jointly sample actions and vision in parallel, by leveraging the bidirectional dependencies in which generated actions constrain plausible future states, and predicted states guide action generation. However, the two modalities are not symmetric in their requirements: image embedding diffusion operates in a high-dimensional space and typically benefits from many denoising steps, whereas low-dimensional action diffusion often converges in far fewer steps. To address this disparity and exploit our decoupled design, we introduce a test-time scaling strategy based on asynchronous forward Euler sampling.
@@ 5 Experiments
VLM backbone and diffusion architecture. We adopt the Eagle-2 model (Li et al., 2025b) as our frozen VLM backbone to process image observations and task instructions. Semantic features are extracted from its 12th layer of the VLM and used as conditioning signals for the diffusion module. The diffusion backbone consists of 12 MMDiT blocks for joint-modal processing, followed by 4 modality-specific DiT blocks for both the action and vision streams. Conditioning with VLM features is applied in an interleaved manner, with alternating self-attention and cross-attention layers.
World modeling target. We follow recent works in avoiding direct pixel-level prediction. The world-modeling target $\tilde{o}_{t+k}$ is the future image embedding derived from the SIGLIP-2 (Tschannen et al., 2025) representations produced by the Eagle-2 model, providing a rich, semantic target for the vision stream to predict. Each image yields 256 tokens from the embedding model, which are reduced to 64 tokens via $2\times 2$ average pooling. In total, the diffusion module process 1 state token, 16 action tokens, and 64 future image tokens. For our joint loss (Eq. 4), we set $\lambda_{WM}=1.0$, equally weighting the action and world-modeling objectives based on our ablation study (Section 5.4).
@@ 5.1 Main results
Table 1: Evaluation on RoboCasa. Success rates (%) on RoboCasa benchmark for 8 pick-and-place (PnP), 6 contraption open/close (OP/CL), and 10 other miscellaneous tasks. 100, 300, and 1,000 demos are used for training. : reproduced results.
@@
Table 2: Evaluation on GR-1. Success rates (%) on GR-1 benchmark for 16 pick-and-place (PnP) and 8 articulated (Art.). 300 and 1,000 demos are used. : reproduced results.
@@
Table 3: Evaluation on real-world tasks. Success rates (%) of 4 pick-and-place (PnP) tasks for real-world Franka robot experiments. See Fig. 4 for the task instructions. : reproduced results.
@@
RoboCasa kitchen. RoboCasa is a single arm manipulation benchmark with a focus on kitchen environment interaction tasks. We utilize a suite of 24 tasks, including turning sink faucets, closing drawer doors, and moving objects. The training dataset is drawn from the publicly available dataset from RoboCasa. We experiment over 100, 300, and 1000 training episodes per tasks.
GR-1 tabletop tasks. GR-1 is a humanoid robot benchmark with a focus on dexterous tabletop manipulation of everyday objects. We utilize a total of 24 tasks, mostly comprised of pick-and-place tasks, with some tasks having additional articulated requirements, such as closing a drawer or microwave. The training dataset is taken from GR00T-N1.5 (NVIDIA et al., 2025). We experiment over 300 and 1000 training episodes per task.
@@
Real-world results. Table 5.1 presents results on Franka robot pick-and-place tasks. DUST consistently outperforms prior methods, achieving the highest success rate on every task with average improvements of 13% over GR00T-N1.5 and 12% over FLARE. These gains, observed across diverse object types and source–target configurations, demonstrate DUST’s robustness in physical environments and its promise for practical deployment. As illustrated in Figure 5, the incorporation of world modeling enables the policy to anticipate the future end-effector pose and accurately align with the target object.
@@ 5.2 Transfer learning
During the pretraining stage, the model is trained exclusively on the video component of the BridgeV2 dataset (Walke et al., 2023), optimizing only the world-modeling term of the flow-matching loss while randomly initializing the action tokens. After pretraining, we finetune the model on the RoboCasa dataset using 100 demonstrations per task. Table 4 shows that incorporating video pretraining yields substantial gains, with DUST achieving an average success rate of 0.585 compared to 0.501 without pretraining. These results highlight that large-scale passive video data can effectively transfer to downstream policy learning, improving data efficiency and generalization while reducing dependence on expensive robot demonstrations.
Figure 5: Qualitative comparison on a real-world pick-and-place task (Instruction : "Pick up the blue cup on the brown box and place it in the golden bowl.") The sequence on the right shows GR00T-N1, which directly generates action sequences, and on the left is DUST, which incorporates explicit world modeling. While GR00T-N1.5 produces actions that bring the gripper near the cup, it fails to align precisely with the rim and is unsuccessful in grasping. By contrast, DUST leverages its internal prediction of future states by estimating where generated actions will position the gripper, allowing it to consistently adjust and achieve alignment with the desired position for grasping.
@@ 5.4 Ablation study
Table 6: Ablation study. Success rates (%) on RoboCasa benchmark with 100 demos/task (a) ablating over architecture and training algorithm, (b) depth of MMDiT, and (c) the loss weigh $\lambda_{\textrm{WM}}$ for world-modeling loss.
@@
Next, we analyze the effect of the loss weighting coefficient $\lambda_{\textrm{WM}}$, which balances the two flow-matching terms in our objective. Larger $\lambda_{\textrm{WM}}$ values emphasize world modeling, while smaller values emphasize action modeling. As shown in Figure 6c, experiments on RoboCasa with 100 demonstrations per task indicate that performance remains stable in the range $\lambda_{\textrm{WM}}\in[0.5,2.0]$, but degrades when moving outside this interval. This suggests that effective learning requires weighting the two objectives relatively evenly. Next, we study the ratio of MMDiT to DiT layers. Fixing the total number of layers in $\pi_{\theta}$ to 16, we vary the number of MMDiT layers to adjust the trade-off between cross-modal knowledge transfer and per-modality specialization. Results (Figure 6b) show that while performance is generally stable across configurations, the best outcome is obtained with 12 MMDiT layers and 4 DiT layers, highlighting the benefit of heavily leveraging cross-modal processing.
@@ 6 Conclusion
In this work, we introduced DUal-STream Diffusion (DUST), a world-model augmented VLA framework that decouples the diffusion of actions and future observations while still enabling cross-modal knowledge transfer. By maintaining separate modality streams linked through shared attention, DUST avoids the limitations of a unified latent space and captures causal dependencies between modalities. Extensive experiments show that DUST consistently outperforms baselines on both simulated benchmarks (RoboCasa, GR-1) and real-world Franka Research 3 tasks, underscoring its scalability and robustness. Beyond architecture and training, we also proposed a test-time scaling strategy with asynchronous joint sampling, which further improves performance by allocating finer-grained diffusion to high-dimensional vision tokens. Finally, pretraining on action-free video (BridgeV2) demonstrates that DUST can exploit large-scale passive data for efficient transfer to downstream robotics. Together, these contributions establish DUST as a versatile and extensible framework for bridging world modeling, video pretraining, and scalable inference in VLA models.
@@ A.1 Test-time scaling of naive joint sampling
In Table 7, we give results to an ablation study, where both $N_{A}=N_{o}$ are increased together, instead of fixing $N_{A}$ and increasing $N_{o}$. We can see that without the decoupling of number of steps between modalities, simply increasing diffusion steps actually leads to deterioration in performance. This lends credibility to our initial hypothesis of only vision tokens needing more diffusion steps, and shows that the asynchronous component of our test-time scaling method is crucial to its success.
@@ A.2 Implementation and Training Details
We base our architecture on the GR00T-N1.5 (NVIDIA et al., 2025) codebase, from which we get the pretrained Eagle-2 VLM model. For vision tokens, they pass through an encoder made up of a 3-layer MLP with 2D sinusoidal positional encoding with SiLU activation. The vision decoder is a 2-layer MLP with ReLU activation. Action tokens utilize the linear encoder-decoder pair given in the original codebase, alongside 1D sinusoidal positional encoding.
@@
Figure 6: Modified MMDiT. MMDiT block is used with separate timestep embeddings being used as conditions for each modality.
@@
The GR00T-N1.5 baseline is trained on the original released code, while the FLARE baseline does not release official code or checkpoints. Hence, for FLARE, we do not utilize the Q-Former architecture of the original paper, but reimplement the FLARE loss to utilize the same world modeling target as ours, which is the SIGLIP embeddings from the model VLM. This allows fair comparison between dual-stream diffusion world modeling of DUST and the implicit world modeling of FLARE. For the alignment module of FLARE we use a small MLP, with similar architecture to that of REPA (Yu et al., 2025), which inspired FLARE.
@@
Excluding batch size and iteration count, all experiments are done with the same training hyperparameters. We optimize with AdamW (Loshchilov & Hutter, 2019) using a base learning rate of $1\text{e-}4$, with $\beta_{1}=0.95$, $\beta_{2}=0.999$, and $\epsilon=1\text{e-}8$. Weight decay of $1\text{e-}5$ is applied with the exemption of bias and LayerNorm weights. The learning rate follows a cosine decay schedule with a 5% warmup period.
@@ A.4 Real-World Experiment Details
Each task contains the four object categories - Teddy Bear, Blue Cube, Blue Cup, and Sponge. During evaluation each object-task configuration gets 6 evaluations, meaning 24 trials per task. We predetermine a set of varied configuration of where to place the source-target locations, on where the source location the object is placed, and the direction it is facing. This allows for more fair comparison in real-world experiments that typically have high stochasticity. When an object has been partially placed in the target destination but the center of gravity is outside of said target, we denote that as a half success and count is as 0.5 successes. We note there were very few cases of this happening.

### v2

@@ 4.3 Vision-action joint sampling and inference-time scaling
During inference, we jointly sample actions and vision in parallel, by leveraging the bidirectional dependencies in which generated actions constrain plausible future states, and predicted states guide action generation. However, the two modalities are not symmetric in their requirements: image embedding diffusion operates in a high-dimensional space and typically benefits from many denoising steps, whereas low-dimensional action diffusion often converges in far fewer steps and even loses performance when sampled over many steps. To address this disparity and exploit our decoupled design, we introduce a test-time scaling strategy based on asynchronous forward Euler sampling.
@@ 5 Experiments
VLM backbone and diffusion architecture. We adopt the Eagle-2 model (Li et al., 2025b) as our frozen VLM backbone to process image observations and task instructions. Semantic features are extracted from the 12th layer of the VLM and used as conditioning signals for the diffusion module. The diffusion backbone consists of 12 MMDiT blocks for joint-modal processing, followed by 4 modality-specific DiT blocks for both the action and vision streams. Conditioning with VLM features is applied in an interleaved manner, with alternating self-attention and cross-attention layers.
World-modeling target. We follow recent works in avoiding direct pixel-level prediction. The world-modeling target $\tilde{o}_{t+k}$ is the future image embedding derived from the SIGLIP-2 (Tschannen et al., 2025) representations produced by the Eagle-2 model, providing a rich, semantic target for the vision stream to predict. Each image yields 256 tokens from the embedding model, which are reduced to 64 tokens via $2\times 2$ average pooling. In total, the diffusion module processes 1 state token, 16 action tokens, and 64 future image tokens. For our joint loss (Eq. 4), we set $\lambda_{WM}=1.0$, equally weighting the action and world-modeling objectives based on our ablation study (Section 5.4).
@@ 5.1 Main results
Table 1: Evaluation on RoboCasa. Success rates (%) on RoboCasa benchmark for 8 pick-and-place (PnP), 6 contraption open/close (OP/CL), and 10 other miscellaneous tasks. 100, 300, and 1,000 demos per task are used for training. : reproduced results.
@@
Table 2: Evaluation on GR-1. Success rates (%) on GR-1 benchmark for 16 pick-and-place (PnP) and 8 articulated (Art.) tasks. 300 and 1,000 demos per task are used. : reproduced results.
@@
Table 3: Evaluation on real-world tasks. Success rates (%) of 4 diverse pick-and-place (PnP) tasks for real-world Franka Research 3 robot experiments. See Fig. 4 for the task instructions and settings. : reproduced results.
@@
RoboCasa kitchen. RoboCasa is a single arm manipulation benchmark with a focus on kitchen environment interaction tasks. We utilize a suite of 24 tasks, including turning sink faucets, closing drawer doors, and moving objects. The training dataset is drawn from the publicly available dataset offered by RoboCasa (Nasiriany et al., 2024). We experiment over 100, 300, and 1000 training episodes per task as training data.
GR-1 tabletop tasks. GR-1 is a humanoid robot benchmark with a focus on dexterous tabletop manipulation of everyday objects. We utilize a total of 24 tasks, mostly comprised of pick-and-place tasks, with some tasks having additional articulated requirements, such as closing a drawer or microwave. The training dataset is taken from GR00T-N1.5 (NVIDIA et al., 2025). We experiment over 300 and 1000 training episodes per task as training data.
@@
Real-world results. Table 5.1 presents results on the Franka Research 3 robot with pick-and-place tasks. DUST consistently outperforms baseline models, achieving the highest success rate on every task with average improvements of 13% over GR00T-N1.5 and 12% over FLARE. These gains, observed across diverse object types and source–target configurations, demonstrate DUST’s robustness in physical environments and its promise for practical deployment. As illustrated in Figure 5, the incorporation of world-modeling enables the policy to anticipate the future end-effector pose and accurately align with the target object.
@@ 5.2 Transfer learning
For this section, during the pretraining stage, the model is trained exclusively on the video component of the BridgeV2 dataset (Walke et al., 2023), optimizing only the world-modeling term of the flow matching loss while randomly initializing the action tokens. After pretraining, we finetune the model on the RoboCasa dataset using 100 demonstrations per task. Table 4 shows that incorporating video pretraining yields substantial gains, with DUST achieving an average success rate of 0.585 compared to 0.501 without pretraining. These results highlight that large-scale passive video data can effectively transfer to downstream policy learning, improving data efficiency and generalization while reducing dependence on expensive robot demonstrations.
Figure 5: Qualitative comparison on a real-world pick-and-place task (Instruction : "Pick up the blue cup on the brown box and place it in the golden bowl.") The sequence on the right shows GR00T-N1, which directly generates action sequences, and on the left is DUST, which incorporates explicit world-modeling. While GR00T-N1.5 produces actions that bring the gripper near the cup, it fails to align precisely with the rim and is unsuccessful in grasping. By contrast, DUST leverages its internal prediction of future states by estimating where generated actions will position the gripper, allowing it to consistently adjust and achieve alignment with the desired position for grasping.
@@ 5.4 Ablation study
Table 6: Ablation study. Success rates (%) on RoboCasa benchmark with 100 demos/task ablating over (a) architecture and training algorithm, (b) depth of MMDiT, and (c) the loss weight $\lambda_{\textrm{WM}}$ for world-modeling loss.
@@
Next, we analyze the effect of the loss weighting coefficient $\lambda_{\textrm{WM}}$, which balances the two flow matching terms in our objective. Larger $\lambda_{\textrm{WM}}$ values emphasize world-modeling, while smaller values emphasize action modeling. As shown in Figure 6c, experiments on RoboCasa with 100 demonstrations per task indicate that performance remains stable in the range $\lambda_{\textrm{WM}}\in[0.5,2.0]$, but degrades when moving outside this interval. This suggests that effective learning requires weighting the two objectives relatively evenly. Next, we study the ratio of MMDiT to DiT layers. Fixing the total number of layers in $\pi_{\theta}$ to 16, we vary the number of MMDiT layers to adjust the trade-off between cross-modal knowledge transfer and per-modality specialization. Results (Figure 6b) show that while performance is generally stable across configurations, the best outcome is obtained with 12 MMDiT layers and 4 DiT layers, highlighting the benefit of heavily leveraging cross-modal processing.
@@ 6 Conclusion
In this work, we introduced dual-stream diffusion (DUST), a world-model augmented VLA framework that decouples the diffusion of actions and future observations while still enabling cross-modal knowledge transfer. By maintaining separate modality streams linked through shared attention, DUST avoids the limitations of a unified latent space and captures causal dependencies between modalities. Extensive experiments show that DUST consistently outperforms baselines on both simulated benchmarks (RoboCasa, GR-1) and real-world Franka Research 3 tasks, underscoring its scalability and robustness. Beyond architecture and training, we also proposed a test-time scaling strategy with asynchronous joint sampling, which further improves performance by allocating finer-grained diffusion to high-dimensional vision tokens. Finally, pretraining on action-free video (BridgeV2) demonstrates that DUST can exploit large-scale passive data for efficient transfer to downstream robotics. Together, these contributions establish DUST as a versatile and extensible framework for bridging world-modeling, video pretraining, and scalable inference in VLA models.
@@ A.1 Test-time scaling of naive joint sampling
In Table 7, we present results from an ablation study, where both $N_{A}=N_{o}$ are increased together, instead of fixing $N_{A}$ and increasing $N_{o}$. We can see that without the decoupling of number of steps between modalities, simply increasing diffusion steps actually leads to deterioration in performance. This lends credibility to our initial hypothesis of only vision tokens needing more diffusion steps, and shows that the asynchronous component of our test-time scaling method is crucial to its success.
@@ A.2 Implementation and Training Details
We base our architecture on the GR00T-N1.5 (NVIDIA et al., 2025) codebase, from which we get the pretrained Eagle-2 VLM model. For vision tokens, they pass through an encoder made up of a 3-layer MLP with 2D sinusoidal positional encoding with SiLU activation. The vision decoder is a 2-layer MLP with ReLU activation. Action tokens utilize the linear encoder-decoder pair given in the original code-base, alongside 1D sinusoidal positional encoding.
@@
Figure 6: Modified MMDiT. DUST’s MMDiT blocks are implemented with separate timestep embeddings being used as conditions for each modality.
@@
The GR00T-N1.5 baseline is trained on the original released code, while the FLARE baseline does not release official code or checkpoints. Hence, for FLARE, we do not utilize the Q-Former architecture of the original paper, but re-implement the FLARE loss to utilize the same world modeling target as ours, which is the SIGLIP-2 embeddings from the model VLM. This allows fair comparison between dual-stream diffusion world modeling of DUST and the implicit world modeling of FLARE. For the alignment module of FLARE we use a small MLP, with similar architecture to that of REPA (Yu et al., 2025), which inspired FLARE.
@@
Excluding batch size and iteration count, all experiments are done with the same training hyperparameters. We optimize with AdamW (Loshchilov & Hutter, 2019) using a base learning rate of $1\text{e-}4$, with $\beta_{1}=0.95$, $\beta_{2}=0.999$, and $\epsilon=1\text{e-}8$. Weight decay of $1\text{e-}5$ is applied with the exception of bias and LayerNorm weights. The learning rate follows a cosine decay schedule with a 5% warmup period.
@@ A.4 Real-World Experiment Details
Each task contains the four object categories - Teddy Bear, Blue Cube, Blue Cup, and Sponge. During evaluation each object-task configuration gets 6 evaluations, meaning 24 trials per task. We predetermine a set of varied configurations of where to place the source-target locations, on where the source location the object is placed, and the direction it is facing. This allows for more fair comparison in real-world experiments that typically have high stochasticity. When an object has been partially placed in the target destination but the center of gravity is outside of said target, we denote that as a half success and count it as 0.5 successes. We note there were very few cases of this happening.